Disk VSN: HOB+0076
Data volume at beginning: 6085.3GB
- 1730 UT: Experiment looked to have started okay but the teslecope drives were stuck. (Arwin)
- 1922 UT: With Brett and Jamie's help got the drives going again from Mt pleasant. First uselful scan is 201-1929 starting at 1929 UT.(Arwin)
- 1954 UT: lots of “ ch -308 ve total power integrator overflow” errors. I presume these are due to the 100 k system temp on channels 14 and 15. (Arwin)
- 202.1327UT: changed to module HOB+0114. Missed scan 202-1324, source 2145+0.82. First source on new module 1736+324 (Lucas).